Lydex Company
Comparative Balance Sheet
This Year Last Year
Assets
Current assets:
Cash $ 960,000 $ 1,200,000
Marketable securities 0 300,000
Accounts receivable, net 2,700,000 1,800,000
Inventory 3,600,000 2,000,000
Prepaid expenses 260,000 200,000
Total current assets 7,520,000 5,500,000
Plant and equipment, net 9,520,000 9,050,000
Total assets $ 17,040,000 $ 14,550,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 4,010,000 $ 2,980,000
Note payable, 10% 3,660,000 3,060,000
Total liabilities 7,670,000 6,040,000
Stockholders' equity:
Common stock, $75 par value 7,500,000 7,500,000
Retained earnings 1,870,000 1,010,000
Total stockholders' equity 9,370,000 8,510,000
Total liabilities and stockholders' equity $ 17,040,000 $ 14,550,000
Lydex Company
Comparative Income Statement and Reconciliation
This Year Last Year
Sales (all on account) $ 15,860,000 $ 13,580,000
Cost of goods sold 12,688,000 10,185,000
Gross margin 3,172,000 3,395,000
Selling and administrative expenses 1,006,000 1,604,000
Net operating income 2,166,000 1,791,000
Interest expense 366,000 306,000
Net income before taxes 1,800,000 1,485,000
Income taxes (30%) 540,000 445,500
Net income 1,260,000 1,039,500
Common dividends 400,000 519,750
Net income retained 860,000 519,750
Beginning retained earnings 1,010,000 490,250
Ending retained earnings $ 1,870,000 $ 1,010,000

To begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex Company’s industry:

Current ratio 2.4
Acid-test ratio 1.1
Average collection period 40 days
Average sale period 60 days
Return on assets 9.3 %
Debt-to-equity ratio 0.7
Times interest earned ratio 5.9
Price-earnings ratio 10

You decide next to assess the company’s stock market performance, Assume that Lydex’s stock price at the end of this year is $94 per share and that at the end of last year it was $62, For both this year and last year, compute: (Round your "Percentage" answers to 1 decimal place and other intermediate and final answers to 2 decimal places)

The earnings per share.

The dividend yield ratio.

The dividend payout ratio.

The price-earnings ratio.

The book value per share of common stock.
Answer #1

1. EPS = Income attributable to common stock holder/number of common stock

= 1,260,000/100000 = 12.60

2. DYR = Dividend per share/average market price per share

= 4/78 = 5.13%

3. DPOR = Dividend per share /EPS

= 4/12.60= 31.75%

4. PER = Average market price per share/EPS

= 78/12.60= 6.19

5. BVPS= Total book value attributable to common stock holder/number of common stock

= 9,370,000/100,000= 93.70